2. Linings of tanks can fail because of salt contamination of the surface that remains after the surface is prepared for the application of the lining. Between 15% and 80% of coating failures have been attributed to residual salt con- tamination. The cost of reworking a failed lining of a specific tank has been estimated at $174,000. [Reference: H. Peters, Monetizing the risk of coating failure, Materials Performance 45(5), 30 (2006).] (a) Calculate the risk due to this type of failure assuming that 20% of failures are caused by residual salt contamination. (b) If the cost of testing and removal of contaminating salts is $4100, is this additional cost justified based on the risk calculation in (a)? (c) Calculate the minimum percentage of failures caused by residual salt contamina- tion at which the additional cost of $4100 for testing and removal of these salts is justified.
Answer #1

(a) risk= prob of event * cost of event = 0.20*174000= $34800

(b) the cost of testing and removal is just $4100 which is much less than 20% probability that failure is caused die to salt contamination. It's hardly 2.35% of the total cost of reworking thus cost of testing and removal is definitely better than option (a).

(c) (4100/17400)*100= 2.35% is the minimum percentage.